OpenStudy (anonymous):

Evaluate f(x) = -x^2 + 1 for x = –3.

OpenStudy (anonymous):

you want to plug in (-3) in place of x. so it will look like the following: (-(-3))^2 + 1 = 9+1 = 10. (Remember that a --3 = a +3

OpenStudy (anonymous):

A. –9 B. –4 C. –8 D. 4

OpenStudy (anonymous):

sorry about that,I got ahead of myself: the anwer is -8. I got it by: -(-3)^2 (you are only squaring the negative 3 first, and then that gives you a 9, but you have a negative in front of that, and that gives you a -9 +1 = -8

OpenStudy (anonymous):

sorry for leading you incorrectly at first.

OpenStudy (anonymous):

its ok thanks for helping

OpenStudy (anonymous):

-x^2 + 1 for x = –3 -> -(-3)^2 +1 = - 9 + 1 = -8